OpenStudy (anonymous):

rewrite cos 3x with only sinx and cosx

OpenStudy (welshfella):

Hint:- cos 3x = cos (2x + x)

OpenStudy (welshfella):

use the identity cos (a + b) = cos a cos b - sin a sin b

OpenStudy (welshfella):

put a = 2x and b = x

OpenStudy (welshfella):

i'll start it off for you cos 3x = cos (2x + x) = cos 2x cos x - can you finish it off?

OpenStudy (anonymous):

\[\cos (A \pm B)=cosAcosB \mp\ sinAsinB\]